OpenStudy (anonymous):

Help Please<3 If an object is dropped from a height of 200 feet, the function h(t)=-16t^2+200 gives the height of the object after t seconds. Approximately, when will the object hit the ground? A. 200.00 sec B. 184.00 Sec C. 3.54 D. 0.78 sec

OpenStudy (gorv):

when object hit the ground then its height should be =f(t)=0

OpenStudy (gorv):

sorry h(t)=0

OpenStudy (anonymous):

Thanks So Much Guys, R U Sure?:)<3 so it would be D?

OpenStudy (gorv):

so 0=-16t^2+200 16t^2=200 t^2=200/16 t=root(200/16)=10*root(2)/4

OpenStudy (anonymous):

Humm thats confusing

OpenStudy (anonymous):

This is embarrassing>.< i hate math i can never get it:(

OpenStudy (gorv):

okkk

OpenStudy (gorv):

0=-16t^2+200

OpenStudy (gorv):

16t^2=200

OpenStudy (gorv):

u got till here ??? @Monica75

OpenStudy (anonymous):

yes

OpenStudy (gorv):

t^2=200/16 t=root(200/16)

OpenStudy (gorv):

t=root(200/16)=3.535=3.54 sec

OpenStudy (gorv):

now ???

OpenStudy (gorv):

@Monica75

OpenStudy (anonymous):

OHHH i see !its all simple steps, thanks for showing me! and taking your time out to help

OpenStudy (gorv):

wlcme :)